A5: Packing Crate Inventory MGT305 You’ve recently been hired as a project manager at a distributor of fruits and vegetables located in Northern Kentucky. An operations manager has asked for your help in analyzing current inventory policies related to packing crates. She’s looking for a lower cost policy. Currently, the operation buys crates from a regional supplier. The operation uses 1800 crates each month. Annual carrying cost is 18% of the $6 acquisition cost per crate. Each order costs approximately $40. Crate supplier lead time is 5 days. Currently, the operation orders crates every two months. You’ve been asked to develop a lower cost alternative to the existing crate inventory policy. Some issues to consider: What is the basis behind the alternative policy that you are proposing and why is it a compelling alternative? Compare the features of each policy (Q, # cycles per year, T, r). Compare the costs of each policy (ordering, holding, TAC). What is a conservative next step (CNS) based on your analysis? A couple of things to consider. Is the reduction in cost of the proposed policy worth the hassle of changing people’s behavior? If cost saving in the case of crates seems low, keep in mind that crates are likely one of many inputs ordered from suppliers. Perhaps trying the new policy on crates, and then expanding the approach to other items if the initial trial is successful, would be worthwhile. Assuming that the crate savings represents the average savings on an item ordered using the new policy approach, and that the operation orders 100 different items, what would be the projected annual savings if the policy was broadly implemented? What are the limitations of this analysis?

Answer the questions below...

Question 15 pts

Inventory policies based on finding the economic order quantity (EOQ) are compelling because they:

Provide an exact amount to order.

Require close coordination with suppliers to implement.

Minimize total costs related to ordering and holding inventory.

Assume that demand is constant.

Permit precise calculation of inventory ordering costs.

Question 25 pts

Currently, the operation orders ____ crates at a time.

21,600

3,600

3,000

6,000

1,800

Question 35 pts

The EOQ for crates is approximately:

365

1,155

2,852

1,265

516

Question 45 pts

The total annual cost of the EOQ policy is about ____ the current policy.

$4200 higher than

about $818 lower than

about $1570 lower than

the same as

about $613 lower than

Question 55 pts

The reorder point for the EOQ policy is approximately:

250 crates

474 crates

296 crates

25 crates

49 crates

Question 65 pts

Compared to the existing policy, the EOQ policy would require ordering less often.

True

False
Answer #1

Annual demand, D = 1800×12 = 21,600 crates

Ordering cost, S = 40

Holding cost, H = 6 × 18% = 1.08

------------------

Inventory policies based on finding the economic order quantity (EOQ) are compelling because they:

Minimize total costs related to ordering and holding inventory.

Currently, the operation orders ____ crates at a time.

3600

Currently, the operation orders every two months. Monthly demand is 1800 crates.

So, order quantity = 1800×2 = 3600

------------------

Question 3. The EOQ for crates is approximately:

1265

Explanation:

EOQ = sqrt(2DS/H)

= sqrt(2×21600×40/1.08)

= 1265 crates

------------------

Question 4. The total annual cost of the EOQ policy is about ____ the current policy.

about $818 lower than

Explanation:

Total annual cost of current policy = Holding cost + Ordering cost =H×Q/2+S×D/Q

= 1.08×3600/2+40×21600/3600

= 2184

Total annual cost of EOQ policy = 1.08×1265/2+40×21600/1265 = 1366

Difference = 2184-1366 = 818

------------------

Question 5. The reorder point for the EOQ policy is approximately:

296 crates

Explanation: Reorder point = Daily demand ×Lead time in days

= (D/365)×5

= (21600/365)×5

= 296 crates,

------------------

Question 6. Compared to the existing policy, the EOQ policy would require ordering less often.

False